# PAT: Pruning-Aware Tuning for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) excel in language tasks, especially with supervised fine-tuning after pre-training. However, their substantial memory and computational requirements hinder practical applications. Structural pruning, which reduces less significant weight dimensions, is one solution. Yet, traditional post-hoc pruning often leads to significant performance loss, with limited recovery from further fine-tuning due to reduced capacity. Since the model fine-tuning refines the general and chaotic knowledge in pre-trained models, we aim to incorporate structural pruning with the fine-tuning, and propose the Pruning-Aware Tuning (PAT) paradigm to eliminate model redundancy while preserving the model performance to the maximum extend. Specifically, we insert the innovative Hybrid Sparsification Modules (HSMs) between the Attention and FFN components to accordingly sparsify the upstream and downstream linear modules. The HSM comprises a lightweight operator and a globally shared trainable mask. The lightweight operator maintains a training overhead comparable to that of LoRA, while the trainable mask unifies the channels to be sparsified, ensuring structural pruning. Additionally, we propose the Identity Loss which decouples the transformation and scaling properties of the HSMs to enhance training robustness. Extensive experiments demonstrate that PAT excels in both performance and efficiency. For example, our Llama2-7b model with a 25% pruning ratio achieves 1.33$\times$ speedup while outperforming the LoRA-finetuned model by up to 1.26% in accuracy with a similar training cost. Code: https://github.com/kriskrisliu/PAT˙Pruning-Aware-Tuning
@@ Introduction
Large language models (LLMs) (Touvron et al. 2023a; Brown et al. 2020; Chowdhery et al. 2022) have transformed the field of NLP (Vaswani et al. 2017; Bahdanau, Cho, and Bengio 2014; Zhang, Zhao, and LeCun 2015; Yang et al. 2016) with their exceptional performance on various complex language benchmarks. Despite their success, these models often necessitate substantial computational resources and present challenges for practical deployment due to their billions of parameters. Their extensive scales result in high latency and complications in deployments (Pan et al. 2023; Zhang et al. 2024). To mitigate these issues, various techniques have been proposed, including model pruning (Ma, Fang, and Wang 2023; Ashkboos et al. 2024; Sun et al. 2023; Santacroce et al. 2023; Fang, Ma, and Wang 2023), knowledge distillation (Agarwal et al. 2023; Tunstall et al. 2023; Sun et al. 2019, 2020; Ma et al. 2020), and quantization (Liu et al. 2022; Yao et al. 2022; Bai et al. 2020; Zafrir et al. 2019) within the context of pre-trained language models (PLMs).
@@
Network pruning (LeCun, Denker, and Solla 1989) has long been recognized as an effective method for model compression and acceleration. Earlier research primarily focused on small-scale networks (Fang et al. 2023; Yang et al. 2023). However, with the advent of large-scale models, pruning techniques have increasingly been applied to large language models (LLMs). According to the pruning granularity, pruning methods can be broadly categorized into unstructured and structured pruning. In the realm of unstructured pruning (Frantar and Alistarh 2023; Sun et al. 2023), techniques such as SparseGPT (Frantar and Alistarh 2023) and Wanda (Sun et al. 2023) have been proposed. SparseGPT addresses the layer-wise reconstruction problem by utilizing Hessian inverses, while Wanda employs the product of weight magnitudes and input feature norms as its pruning criterion. Despite their effectiveness, these unstructured sparsification methods do not guarantee on-device speedup without hardware-specific support. In contrast, the structured pruning (Zafrir et al. 2021; Kurtic et al. 2022; Xia, Zhong, and Chen 2022; Yang, Wen, and Li 2019; Yang et al. 2023) removes organized patterns within the network, enabling significant acceleration in a hardware-agnostic manner. For instance, Shortened-LLaMA (Kim et al. 2024) introduces a block pruning method to remove Transformer blocks, resulting in depth pruning. Sheared-LLaMA (Xia et al. 2023) incorporates a mask learning phase to identify prunable components across both the network’s width and depth. LLM-Pruner (Ma, Fang, and Wang 2023) and SliceGPT (Ashkboos et al. 2024) eliminates coupled structures in the aspect of network width while retaining the number of layers: LLM-Pruner sparsifies the intermediate dimension while SliceGPT focuses on the hidden dimension. However, many existing structured pruning models still suffer from accuracy loss, necessitating further exploration and improvement.
@@
We employ the LaMini-instruction dataset (Wu et al. 2023) for fine-tuning. To reduce training costs, we randomly drop 50% of the samples, resulting in a final dataset of 1 million samples. Unless otherwise stated, all experimental results are based on this setting. We conduct zero-shot evaluation on 14 datasets, including ARC-Challenge (Clark et al. 2018), ARC-Easy (Clark et al. 2018), BOOLQ (Wang et al. 2019a), COPA (Wang et al. 2019a), HellaSwag (Zellers et al. 2019), MMLU (Hendrycks et al. 2021), MultiRC (Wang et al. 2019a), OpenBookQA (Mihaylov et al. 2018), PIQA (Bisk et al. 2020), RTE (Wang et al. 2019a), SIQA (Sap et al. 2019), WIC (Wang et al. 2019a), WinoGrande (ai2 2019), WSC (Wang et al. 2019a). The accuracy is calulated by First-Capital-Word [^2] https://github.com/open-compass/opencompass (Contributors 2023) method.
@@
Experiments are conducted using A100 GPUs. The models are fine-tuned over 3 epochs using the Alpaca instruction template. The learning rate is set to $5\times 10^{-5}$ with a cosine schedule. The batch size is set to 128, and the sequence length is 256 tokens. The milestone step of our PAT, $s_{0}$, is set to $1/3$ of the total training steps. The settings of our HIOs are derived to match the number of trainable parameters with LoRA-64, as detailed in the Appendix.
@@
The implementation of HIO significantly reduces the number of trainable parameters, but this reduction may directly impact the model accuracy. We conducted experiments using various scales of training parameters on the Llama 2 7B model, and illustrate the results in Fig. 4. The total number of trainable parameters is adjusted by the rank values of HIO and LoRA modules. For example, the “LoRA-64” in dark represents the traditional LoRA fine-tuning with a rank value set to 64, and the “HIO-200, LoRA-20” in purple represents our PAT with a rank of 200 in HIO and a rank of 20 in LoRA modules. We find that our PAT demonstrates a performance trend correlated to the number of trainable parameters. “Dense [^3] Indicating that we use the dense matrix instead of the HIO., LoRA-8” with 14.15% trainable parameters achieves 64.19% accuracy, outperforming “LoRA-64” by 5.43%. Conversely, “HIO-8, LoRA-8” with merely 0.36% trainable parameters results in a 6% accuracy reduction. In practice, we opt for “HIO-200, LoRA-20” in Llama 2 7B experiments, aligning the parameter count with that of “LoRA-64”. Details for settings on other models are provided in the Appendix.
@@
Table 2: Ablation study on trainable mask and identity loss.
@@
Following the downstream task adaptation detailed in DoRA (Liu et al. 2024a), we leverage PAT to fine-tune on specific tasks, including ARC, SuperGlue, OpenBookQA, PIQA, SIQA, MMLU, and WinoGrande, totaling 387,527 samples. The setting of HSMs is “HIO-200, LoRA/DoRA-20”. Our 25% pruned PAT-L and PAT-D achieve performance levels on par with those achieved by traditional DoRA and LoRA.
@@ Conclusion
In this paper, we propose Pruning-Aware Tuning (PAT), a novel structured pruning approach for Large Language Models (LLMs). PAT focuses on sparsifying the hidden dimensions during the fine-tuning stage, while preserving the linguistic capabilities of LLMs, even under significant pruning ratios. We utilize a single trainable mask to adaptively set channel values of hidden states to zero. We develop efficient Hybrid Sparsification Modules (HSMs) that are inserted between the Attention and FFN blocks, enabling pruning of all linear layers accordingly. It reduces PAT’s training overhead to levels comparable to traditional LoRA fine-tuning. Additionally, we propose the Identity Loss to enhance the training robustness by decoupling the rotation and scaling properties of the HSMs. In the zero-shot evaluation, our 30%-PAT Llama2 7B and 13B models maintains 98% performance of those achieved from the LoRA fine-tuning.
Appendix
This section serves as supplementary materials to the main paper. We give details on the settings of Hybrid Identity Operator. We also provide detailed results of zero-shot evaluation.
HIO Settings
We meticulously select our PAT settings in alignment with the corresponding LoRA configurations. The specific settings are detailed in Tab. 4. For example, in the case of Gemma 2B, when the traditional LoRA method employs a rank value of 64, our PAT approach adjusts this to a rank value of 20 for LoRA and 300 for HIO. This configuration results in a smaller proportion of trainable parameters in PAT compared to LoRA, while still maintaining model accuracy even at high pruning ratios. We test several HIO settings in the Llama2 7B model to discuss the efficiency and accuracy trade-off. We find that the small number of trainable parameters in HIO will lead to poor accuracy. In practice, we choose a rank value of 200 for HIO in Llama2 and Yi-1.5 models, and 300 for that in Gemma models. The rank value for LoRA can be therefore determined accordingly.
@@
Main Results
The zero-shot evaluation results of LoRA (Hu et al. 2021), LLM-Pruner (Ma, Fang, and Wang 2023), SliceGPT (Ashkboos et al. 2024) and our PAT are shown in Tabs. 5, 6, 9, 8 and 7. The evaluation tasks consist of ARC-C (Clark et al. 2018), ARC-E (Clark et al. 2018), BOOLQ (Wang et al. 2019a), COPA (Wang et al. 2019a), HS (Zellers et al. 2019), MMLU (Hendrycks et al. 2021), MULTIRC (Wang et al. 2019a), OBQA (Mihaylov et al. 2018), PIQA (Bisk et al. 2020), RTE (Wang et al. 2019a), SIQA (Sap et al. 2019), WIC (Wang et al. 2019a), WG (ai2 2019), and WSC (Wang et al. 2019a). We highlight our results in the orange background.
Llama2
We employ Llama2 7B and 13B models. As demonstrated in Tabs. 5 and 6, the Llama2 models exhibit a higher tolerance to pruning compared to other models. Notably, at a 30% pruning ratio, our PAT method retains 98% of the zero-shot accuracy when compared to the LoRA-64 approach. Interestingly, our PAT method at 20% and 25% pruning ratios even outperforms LoRA-64.
Gemma
We employ Gemma 2B and 7B models. The Gemma 2B encounters significant accuracy degradation when reaches a pruning ratio of 30%. The Gemma 7B exhibits linear degradation according to pruning ratios.

### v2

@@ Abstract
Large language models (LLMs) excel in language tasks, especially with supervised fine-tuning after pre-training. However, their substantial memory and computational requirements hinder practical applications. Structural pruning, which reduces less significant weight dimensions, is one solution. Yet, traditional post-hoc pruning often leads to significant performance loss, with limited recovery from further fine-tuning due to reduced capacity. Since the model fine-tuning refines the general and chaotic knowledge in pre-trained models, we aim to incorporate structural pruning with the fine-tuning, and propose the Pruning-Aware Tuning (PAT) paradigm to eliminate model redundancy while preserving the model performance to the maximum extend. Specifically, we insert the innovative Hybrid Sparsification Modules (HSMs) between the Attention and FFN components to accordingly sparsify the upstream and downstream linear modules. The HSM comprises a lightweight operator and a globally shared trainable mask. The lightweight operator maintains a training overhead comparable to that of LoRA, while the trainable mask unifies the channels to be sparsified, ensuring structural pruning. Additionally, we propose the Identity Loss which decouples the transformation and scaling properties of the HSMs to enhance training robustness. Extensive experiments demonstrate that PAT excels in both performance and efficiency. For example, our Llama2-7b model with a 25% pruning ratio achieves 1.33$\times$ speedup while outperforming the LoRA-finetuned model by up to 1.26% in accuracy with a similar training cost.
Codehttps://github.com/kriskrisliu/PAT
@@ Introduction
Large language models (LLMs) (Touvron et al. 2023a; Brown et al. 2020; Chowdhery et al. 2022) have transformed the field of NLP (Vaswani et al. 2017; Bahdanau, Cho, and Bengio 2014; Zhang, Zhao, and LeCun 2015; Yang et al. 2016) with their exceptional performance on various complex language benchmarks. Despite their success, these models often necessitate substantial computational resources and present challenges for practical deployment due to their billions of parameters. Their extensive scales result in high latency and complications in deployments (Pan et al. 2023; Zhang et al. 2024). To mitigate these issues, various techniques have been proposed, including model pruning (Ma, Fang, and Wang 2023; Ashkboos et al. 2024; Sun et al. 2023; Santacroce et al. 2023; Fang, Ma, and Wang 2023), knowledge distillation (Agarwal et al. 2024; Tunstall et al. 2023; Sun et al. 2019, 2020; Ma et al. 2020), and quantization (Liu et al. 2022; Yao et al. 2022; Bai et al. 2020; Zafrir et al. 2019) within the context of pre-trained language models (PLMs).
@@
Network pruning (LeCun, Denker, and Solla 1989) has long been recognized as an effective method for model compression and acceleration. Earlier research primarily focused on small-scale networks (Fang et al. 2023; Yang et al. 2023; Chen et al. 2021; Wu et al. 2024). However, with the advent of large-scale models, pruning techniques have increasingly been applied to large language models (LLMs). According to the pruning granularity, pruning methods can be broadly categorized into unstructured and structured pruning. In the realm of unstructured pruning (Frantar and Alistarh 2023; Sun et al. 2023), techniques such as SparseGPT (Frantar and Alistarh 2023) and Wanda (Sun et al. 2023) have been proposed. SparseGPT addresses the layer-wise reconstruction problem by utilizing Hessian inverses, while Wanda employs the product of weight magnitudes and input feature norms as its pruning criterion. Despite their effectiveness, these unstructured sparsification methods do not guarantee on-device speedup without hardware-specific support. In contrast, the structured pruning (Zafrir et al. 2021; Kurtic et al. 2022; Xia, Zhong, and Chen 2022; Yang, Wen, and Li 2019; Yang et al. 2023) removes organized patterns within the network, enabling significant acceleration in a hardware-agnostic manner. For instance, Shortened-LLaMA (Kim et al. 2024) removes Transformer blocks, resulting in depth pruning. Sheared-LLaMA (Xia et al. 2023) incorporates the learnable mask to prune both the network’s width and depth. LLM-Pruner (Ma, Fang, and Wang 2023) and SliceGPT (Ashkboos et al. 2024) prune the network width while retaining the number of layers: LLM-Pruner sparsifies the intermediate dimension while SliceGPT focuses on the hidden dimension. However, existing structured pruning models still suffer from accuracy loss, necessitating further exploration and improvement.
@@
We employ the LaMini-instruction dataset (Wu et al. 2023) for fine-tuning. To reduce training costs, we randomly drop 50% of the samples, resulting in a final dataset of 1 million samples. Unless otherwise stated, all experimental results are based on this setting. We conduct zero-shot evaluation on 14 datasets, including ARC-Challenge (Clark et al. 2018), ARC-Easy (Clark et al. 2018), BOOLQ (Wang et al. 2019a), COPA (Wang et al. 2019a), HellaSwag (Zellers et al. 2019), MMLU (Hendrycks et al. 2021), MultiRC (Wang et al. 2019a), OpenBookQA (Mihaylov et al. 2018), PIQA (Bisk et al. 2020), RTE (Wang et al. 2019a), SIQA (Sap et al. 2019), WIC (Wang et al. 2019a), WinoGrande (Sakaguchi et al. 2021), WSC (Wang et al. 2019a). The accuracy is calulated by First-Capital-Word [^3] https://github.com/open-compass/opencompass (Contributors 2023) method.
@@
Experiments are conducted using A100 GPUs. The models are fine-tuned over 3 epochs using the Alpaca instruction template. The learning rate is set to $5\times 10^{-5}$ with a cosine schedule. The batch size is set to 128, and the sequence length is 256 tokens. The milestone step of our PAT, $s_{0}$, is set to $1/3$ of the total training steps. The settings of our HIOs are derived to match the number of trainable parameters with LoRA-64. For example, we set the rank values of HIO and LoRA modules to 200 and 20 in the Llama2-7B experiments, respectively.
@@
The implementation of HIO significantly reduces the number of trainable parameters, but this reduction may directly impact the model accuracy. We conducted experiments using various scales of training parameters on the Llama 2 7B model, and illustrate the results in Fig. 4. The total number of trainable parameters is adjusted by the rank values of HIO and LoRA modules. For example, the “LoRA-64” in dark represents the traditional LoRA fine-tuning with a rank value set to 64, and the “HIO-200, LoRA-20” in purple represents our PAT with a rank of 200 in HIO and a rank of 20 in LoRA modules. We find that our PAT demonstrates a performance trend correlated to the number of trainable parameters. “Dense [^4] Indicating that we use the dense matrix instead of the HIO., LoRA-8” with 14.15% trainable parameters achieves 64.19% accuracy, outperforming “LoRA-64” by 5.43%. Conversely, “HIO-8, LoRA-8” with merely 0.36% trainable parameters results in a 6% accuracy reduction. In practice, we opt for “HIO-200, LoRA-20” in Llama 2 7B experiments, aligning the parameter count with that of “LoRA-64”. For others, Gemma 2B with “HIO-300, LoRA20”, Gemma 7B with “HIO-300, LoRA20”, Llama2 13B with “HIO-200, LoRA20” , and Yi-1.5 34B with “HIO-200, LoRA20”.
@@
Table 2: Ablation study on trainable mask and identity loss.
@@
Following the downstream task adaptation detailed in DoRA (Liu et al. 2024a), we leverage PAT to fine-tune on specific tasks, including ARC, SuperGlue, OpenBookQA, PIQA, SIQA, MMLU, and WinoGrande. The setting of HSMs is “HIO-200, LoRA/DoRA-20”. Our 25% pruned PAT-L and PAT-D achieve performance levels on par with those achieved by traditional DoRA and LoRA, shown in Tab. 3.
@@ Conclusion
We propose Pruning-Aware Tuning (PAT), a novel structured pruning approach for Large Language Models (LLMs). PAT prunes the hidden dimensions during the fine-tuning, while preserving the linguistic capabilities. We develop a trainable mask to adaptively set channel values to zero, and efficient Hybrid Sparsification Modules to enable pruning of all linear layers accordingly. The efficiency design reduces the training overhead of PAT to levels comparable to traditional LoRA fine-tuning. Additionally, we propose the Identity Loss to enhance the training robustness by decoupling the rotation and scaling properties of the HSMs. In the zero-shot evaluation, our 30%-PAT Llama2 7B and 13B models maintains 98% performance of those achieved from the LoRA fine-tuning.
Acknowledgments
This work was supported in part by the Strategic Industries and Key Technologies Project of Jiangsu Province under Grant BE2023020-3.
Appendix A Appendix
This section serves as the appendix to the main paper.
Detailed Main Results
We evaluate LoRA (Hu et al. 2021), LLM-Pruner (Ma, Fang, and Wang 2023), SliceGPT (Ashkboos et al. 2024) and our PAT on 14 tasks. Results are shown in Tabs. 5, 6, 7, 8 and 9.
@@ Appendix A Appendix
Detailed HIO Settings
We present the detailed settings for HIO in Tab. 4. Our experiments are conducted on the LLaMA 2-7B model, varying the rank values in the LoRA module from 8 to 512 and in the HIO module from 8 to 1024. Across all configurations, our PAT method demonstrates consistently strong performance. To strike a balance between the number of trainable parameters and computational overhead, we select the “LoRA20, HIO200” configuration for the main experiments.